I need help explaining this equation to my 12yr old daughter.I need to be able to explain the steps to her. Can you guys help?
8(4+9a)

Found 2 solutions by ewatrrr, JoelSchwartz:
Answer by ewatrrr(24785)

The Step is Distributing the 8 over the Parenthesis:
8(4+ 9a) = 32 + 72a

Answer by JoelSchwartz(130)

When a number is right next to the parentheses it can be distributed. When distributing the eight it can be multiplied by the four and the 9a. The result would be 8+8+8+8+9a+9a+9a+9a+9a+9a+9a+9a=32+72a